Calculate the amount of water (in grams) that must be added to each of the following substances to produce the indicated solution
20.7 g of MgCl2 in the preparation of a 6.2% mass solution
Answer #1

The mass of MgCl2 in the solution should be 6.2% of total mass of the solution. Let the mass of water added be x grams

20.7/(20.7+x) *100 = 6.2

(20.7 +x) = 20.7/6.2 *100

20.7 + x = 333.87

x = 333.87-20.7 = 313.17 g

Answer: 313.17 g